# Uncovering Overfitting in Large Language Model Editing

### v1

@@ Appendix A Details on the EVOKE Benchmark
The EVOKE benchmark is designed to evaluate the effectiveness of knowledge editing methods for LLMs, focusing not only on their ability to accurately recall edited knowledge but also on their performance and potential overfitting across diverse complex reasoning tasks. Notably, part of EVOKE’s data is sourced from existing knowledge editing datasets. We have reorganized and extended datasets including CounterFact (Meng et al., 2022a), CounterFactPlus (Yao et al., 2023), and RippleEdits-Popular (Cohen et al., 2023) to construct new task-specific data for comprehensive analysis of knowledge editing methods.
@@
Appendix B Experimental Setup Details
Our experiments build on the codebase implemented by Meng et al. (2022a; b). For the hyperparameters of MEMIT on GPT-2 XL, we observe significant failures in certain editing cases, with notably lower performance metrics on the recall task compared to GPT-J. To address this, we adjust the hyperparameters, specifically increasing the clamp_norm_factor from $0.75$ to $1.25$ for GPT-2 XL. This adjustment is made to align its editing success rates with those of GPT-J, allowing us to evaluate normal editing performance and potential overfitting. All other baseline implementations, including hyperparameters, remain consistent with the setup of Meng et al. (2022a; b).
B.1 Baseline Methods
@@
B.2 Details on Multi-layer Editing Experiment
In Appendix E, we investigate the impact of distributing parameter updates across multiple layers on Editing Overfit, conducting experiments using MEMIT method. The original MEMIT approach, when applied to GPT-J, edited six layers ($l={3,4,5,6,7,8}$), with the recall vector $\mathbf{v}$ computed at layer $8$. In our experiments, we maintain a constant product of the hyperparameter clamp_norm_factor and the number of edited layers. When adjusting the number of layers, we keep the highest edited layer fixed. For example, when editing three layers, we select $l=\{6,7,8\}$.
B.3 Details on Data Augmentation Experiment
In Section 5, we follow existing data augmentation strategy (Wei et al., 2024; Gangadhar & Stratos, 2024a) to evaluate the effect of data augmentation on Editing Overfit. Specifically, we employ GPT-4o to generate $10$ data-augmented paraphrase prompts and $10$ subject specificity prompts for each edit request. Paraphrases are required to be factual statements semantically equivalent to the edited knowledge, while subject specificity prompts are factual statements about the same subject but with different relations. The construction of specificity augmentation prompts aligned with the task test prompts for Subject Specificity, directly enhancing this task. We also utilize GPT-4o to verify and filter out augmented samples that do not meet the requirements.
@@ B.3 Details on Data Augmentation Experiment
Appendix C Method Implementation Details
@@
The choice of layers for applying the Subject Representation Constraint (SRC) in LTI is a critical hyperparameter, guided by previous interpretability studies on LLM knowledge recall mechanisms(Meng et al., 2022a; Cohen et al., 2023). These analyses reveal that subject position representations, which encode relevant knowledge, are extracted by attention modules in deeper layers for answer prediction. The SRC aims to align the subject representations of the edited model with those produced by normal inference, ensuring that subsequent recall steps function as expected. Specifically, Cohen et al. (2023) highlight that for GPT-2 XL, the attention edges from “subject-to-last-token” in layers $30-40$ are strongly correlated with answer prediction, while for GPT-J, this correlation is observed in layers $15-20$. Therefore, we constrain layer $30$ in GPT-2 XL and layer $15$ in GPT-J, anticipating these representations to be effectively utilized by these information flows. Further experimental results on the effects of constraining different layers are presented in Appendix F.
@@
LTI’s hyperparameters include coefficients for the three constraint losses. In practice, the coefficient $\lambda$ for the Subject Representation Constraint is set to $0.0625$, while the Output Distribution Constraint coefficient $\beta$ is $0.0325$. For the New Knowledge Constraint coefficient $\alpha$, ROME-LTI uses values of $0.0625$ for GPT-J and and $0.15$ for GPT-2 XL, whereas MEMIT-LTI employs $0.25$ for GPT-J and $0.125$ for GPT-2 XL.
Appendix D Rank-One Model Editing
@@
Appendix E Analysis on Distributing Weight Updates Across Layers
@@
Appendix F The impact of Constraint Layers
@@
Figure 9 demonstrates the model’s performance when varying the number of intermediate layers constrained in ROME-LTI. Despite minor fluctuations across different metrics, the performance on all overfit tasks consistently remains significantly better than that of the original ROME method. This suggests that the model’s performance is not highly sensitive to the specific layer position of intermediate constraints. This phenomenon may be attributed to that aligning the hidden state of a specific layer with the target representation likely influences the reasoning process in adjacent layers for the same token position. This implicitly extends the constraint effect to other layers, encouraging them to also approximate the corresponding representations in the target inference process. Previous ablation studies (§6.3) have shown that removing hidden state constraints at the subject position significantly degrades performance, and these results indicate that constraining specific token positions in intermediate layers may be more crucial than constraining particular layer numbers.
Appendix G Case Study

### v2

@@ Appendix C Details on the EVOKE Benchmark
The EVOKE benchmark is designed to evaluate the effectiveness of knowledge editing methods for LLMs, focusing not only on their ability to accurately recall edited knowledge but also on their performance and potential overfitting across diverse complex reasoning tasks. Notably, part of EVOKE’s data is sourced from existing knowledge editing datasets. We have reorganized and extended datasets including CounterFact (Meng et al., 2022a), CounterFactPlus (Yao et al., 2023), and RippleEdits-Popular (Cohen et al., 2024) to construct new task-specific data for comprehensive analysis of knowledge editing methods.
@@
Appendix D Experimental Setup Details
Our experiments build on the codebase implemented by Meng et al. (2022a; b). For the hyperparameters of MEMIT on GPT-2 XL, we observe significant failures in certain editing cases, with notably lower performance metrics on the recall task compared to GPT-J. To address this, we adjust the hyperparameters, specifically increasing the clamp_norm_factor from $0.75$ to $1.25$ for GPT-2 XL. This adjustment is made to align its editing success rates with those of GPT-J, allowing us to evaluate normal editing performance and potential overfitting. All other baseline implementations, including hyperparameters, remain consistent with the setup of Meng et al. (2022a; b), and hyperparameters on Llama-2-7B remain consistent with Yao et al. (2023).
D.1 Baseline Methods
@@
D.2 Details on Multi-layer Editing Experiment
In Appendix H, we investigate the impact of distributing parameter updates across multiple layers on Editing Overfit, conducting experiments using MEMIT method. The original MEMIT approach, when applied to GPT-J, edited six layers ($l={3,4,5,6,7,8}$), with the recall vector $\mathbf{v}$ computed at layer $8$. In our experiments, we maintain a constant product of the hyperparameter clamp_norm_factor and the number of edited layers. When adjusting the number of layers, we keep the highest edited layer fixed. For example, when editing three layers, we select $l=\{6,7,8\}$.
D.3 Details on Data Augmentation Experiment
In Section 5, we follow existing data augmentation strategy (Wei et al., 2024; Gangadhar & Stratos, 2024) to evaluate the effect of data augmentation on Editing Overfit. Specifically, we employ GPT-4o to generate $10$ data-augmented paraphrase prompts and $10$ subject specificity prompts for each edit request. Paraphrases are required to be factual statements semantically equivalent to the edited knowledge, while subject specificity prompts are factual statements about the same subject but with different relations. The construction of specificity augmentation prompts aligned with the task test prompts for Subject Specificity, directly enhancing this task. We also utilize GPT-4o to verify and filter out augmented samples that do not meet the requirements.
@@ D.3 Details on Data Augmentation Experiment
Appendix E Method Implementation Details
@@
The choice of layers for applying the Subject Representation Constraint (SRC) in LTI is a critical hyperparameter, guided by previous interpretability studies on LLM knowledge recall mechanisms(Meng et al., 2022a; Cohen et al., 2024). These analyses reveal that subject position representations, which encode relevant knowledge, are extracted by attention modules in deeper layers for answer prediction. The SRC aims to align the subject representations of the edited model with those produced by normal inference, ensuring that subsequent recall steps function as expected. Specifically, Cohen et al. (2024) highlight that for GPT-2 XL, the attention edges from “subject-to-last-token” in layers $30-40$ are strongly correlated with answer prediction, while for GPT-J, this correlation is observed in layers $15-20$. Therefore, we constrain layer $30$ in GPT-2 XL and layer $15$ in GPT-J and Llama-2-7B, anticipating these representations to be effectively utilized by these information flows. Further experimental results on the effects of constraining different layers are presented in Appendix I.
@@
LTI’s hyperparameters include coefficients for the three constraint losses. In practice, the coefficient $\lambda$ for the Subject Representation Constraint is set to $0.0625$, while the Output Distribution Constraint coefficient $\beta$ is $0.0325$. For the New Knowledge Constraint coefficient $\alpha$, ROME-LTI uses values of $0.0625$ for GPT-J, $0.15$ for GPT-2 XL and $0.35$ for Llama-2-7B, whereas MEMIT-LTI employs $0.25$ for GPT-J and $0.125$ for GPT-2 XL and Llama-2-7B.
Appendix F Rank-One Model Editing
@@
Appendix G Results on Llama-2-7B
In this section, we extend our experiments to the Llama-2-7B (Touvron et al., 2023) model using the EVOKE dataset, with results presented in Tables 5 and 6. Consistent with findings on GPT-2 XL and GPT-J, overfitting patterns persist on this more advanced model: FT continues to overfit to
Table 5: Results (AMS↑ (%)) on the Recall Tasks of EVOKE.
Editor Llama-2-7B
Efficacy Paraphrase
BASE 13.09 15.08
FT 99.61 92.48
FT-L 92.73 21.53
ROME 100.00 93.60
ROME-LTI 100.00 89.42
MEMIT 100.00 96.80
MEMIT-LTI 100.00 91.71
$p(s,r)\rightarrow o*$ pattern, while locate-then-edit methods maintain their overfit pattern to single $s\rightarrow o*$ mappings. These results further confirm the universality of the Editing Overfit phenomenon. Specifically, nearly all successfully edited models exhibit significantly higher Direct Probability (DP) scores compared to the unedited model. The average DP for FT, ROME, and MEMIT on most overfit tasks significantly surpasses the Correct Answer Probability (CAP), with elevated EOS values indicating persistent issues across edited samples. Notably, both ROME-LTI and MEMIT-LTI show significant improvements in overfitting metrics (DP and EOS) compared to their base versions, demonstrating effective mitigation of Editing Overfit.
Editor Prefix Distraction Multi-hop Reasoning Subject Specificity Relation Specificity DP↓ CAP↑ EOS↑ DP↓ OAP↓ CAP↑ EOS↑ AMS↑ DP↓ CAP↑ EOS↑ DP↓ CAP↑ EOS↑ Llama-2-7B 16.42 28.03 58.44 1.80 21.79 9.30 86.01 28.95 0.74 42.37 97.38 1.18 21.21 82.92 FT 51.48 6.11 5.61 27.71 9.26 12.21 35.40 56.20 9.74 11.03 46.07 23.12 6.30 18.83 FT-L 16.90 26.29 53.17 2.59 19.47 10.96 85.16 36.13 1.12 31.62 95.20 1.54 20.14 81.08 ROME 22.80 25.22 46.66 21.66 9.06 24.98 63.99 77.37 21.82 15.23 45.63 2.09 20.52 81.11 ROME-LTI 17.90 27.60 53.13 15.55 11.79 23.48 70.32 71.53 14.30 18.65 57.42 1.73 20.63 81.36 MEMIT 40.76 16.90 25.36 30.47 6.05 22.29 48.18 81.75 39.09 8.93 23.36 3.87 19.47 78.08 MEMIT-LTI 26.75 20.77 36.81 14.15 10.32 21.31 67.52 74.60 9.06 10.86 47.60 2.15 19.74 80.30
Appendix H Analysis on Distributing Weight Updates Across Layers
@@
Appendix I The impact of Constraint Layers
Figure 9 demonstrates the model’s performance when varying the number of intermediate layers constrained in ROME-LTI. Despite minor fluctuations across different metrics, the performance on all overfit tasks consistently remains significantly better than that of the original ROME method. This suggests that the model’s performance is not highly sensitive to the specific layer position of intermediate constraints. This phenomenon may be attributed to that aligning the hidden state of a specific layer with the target representation likely influences the reasoning process in adjacent layers for the same token position. This implicitly extends the constraint effect to other layers, encouraging them to also approximate the
@@ Appendix I The impact of Constraint Layers
corresponding representations in the target inference process. Previous ablation studies (§6.3) have shown that removing hidden state constraints at the subject position significantly degrades performance, and these results indicate that constraining specific token positions in intermediate layers may be more crucial than constraining particular layer numbers.
Appendix J Case Study